# Pseudo Transitions in the Finite-Size Blume-Capel Model

### v2

@@ 3.2 Third order transitions
In Figs. 6(a) and 7(a), the ’backbending’ phenomenon is observed in the first derivative of entropy, $\beta(E)$. By analyzing Figs. 6(b)-(c) and 7(b)-(c), we found that the third-order independent transition occurs in the ’backbending’ region at lower energy levels, whereas the phase transition point occurs in the ’backbending’ region at higher energy levels. Consequently, although their temperatures are very close, they exist in completely different states, with a noticeable energy level difference. By comparing Figs. 6(a) and 7(a), it can be seen that as the value of $D$ increases, the ’backbending’ effect in $\beta(E)$ intensifies. As a result, the positions of the third-order independent transition and the phase transition point to become increasingly close to each other. This phenomenon is also observed in the Potts model, suggesting that it is characteristic of first-order phase transitions. The ’backbending’ phenomenon is a characteristic signal of phase coexistence in first-order phase transitions. The maximum error between the positions of the third-order independent transition and the first-order phase transition point, obtained by both methods, is 0.002, which is negligible.
@@
In Fig. 9, we present the temperature-dependent variations of the average perimeter and its first derivative across different parameters. Our primary focus is to analyze the behavior of the first derivative. In the figure, red vertical lines mark the positions of the critical points, while purple lines indicate the locations of the identified third-order dependent transitions. Corresponding temperature values are annotated next to these lines in matching colors. When compared with the results obtained from microcanonical inflection point analysis, we find that the discrepancies between the two methods are approximately 0.03. Additionally, in Fig. 9(g), we observe a notable distinction, where a weak signal of a third-order dependent transition is detected and marked by a purple vertical line. Through careful analysis of this signal, we observed that the local minimum shifts upward as the system size increases. Consequently, we infer that, in the limit of infinite system size, this signal vanishes, aligning with the results from our microcanonical analysis. In Fig. 9(h), for $D=1.97$, the system has fully entered the first-order phase transition region, we observe that this signal disappears completely. This phenomenon further indicates that in the first-order phase transition region, the signal of a third-order dependent transition may not exist.
@@
Using two different methods, we were able to identify signals of both third-order independent and dependent transitions within the second-order phase transition regime, while only third-order independent transition signals were identified in the first-order phase transition regime. Only at the tricritical point, a very weak signal of the third-order dependent transition was detected through geometric analysis.
@@ 4 Summary
In this paper, we employed the Wang-Landau (WL) sampling method to obtain the density of states (DOS) for the Blume-Capel model, while using the Metropolis sampling method to simulate the spin configurations and extract information on isolated spins and the average cluster perimeter. We performed both canonical and microcanonical ensemble analyses on the data extracted from these two methods, successfully determining the positions of the critical and first-order phase transition points. The results from both sampling methods showed only minor differences, providing strong evidence for the validity and accuracy of our findings.
Importantly, our study revealed an intriguing phenomenon. Through microcanonical inflection point analysis, we observed that both third-order dependent and independent transitions coexist at the critical phase transition, while only third-order independent transitions were observed in the first-order phase transition. This observation was further verified through geometric analysis of the behavior of isolated spins and average cluster perimeters. Similar conclusions have been drawn in the Ising model [18] and the Potts model [19]. Building on these findings, we further hypothesize that third-order dependent transitions may only occur during critical phase transitions and are absent in first-order transitions. This hypothesis provides significant theoretical support for our research on critical early-warning systems and offers valuable guidance for future studies on phase transitions.
Our next step is to extend this methodology to practical liquid systems or more complex magnetic systems, aiming to confirm this hypothesis.
@@ 5 Acknowledgments
We would to thank Professor Ying Tang for his valuable discussions. This work is supported by the National Natural Science Foundation of China NO.12304257.

### v3

@@ 3.2 Higher order transitions
By comparing Tables 2 and 3, we observe that when the primary phase transition of the system is of the first order, the positions of the third-order independent transition and the phase transition point, as determined by microcanonical inflection point analysis, are closely aligned. In Figs. 6(a) and 7(a), the “backbending” phenomenon is observed in the first derivative of entropy, $\beta(E)$. By analyzing Figs. 6(b)-(c) and 7(b)-(c), we found that the third-order independent transition occurs in the “backbending” region at lower energy levels, whereas the phase transition point occurs in the “backbending” region at higher energy levels. Consequently, the temperatures of the third-order independent transition and the phase transition point are very close; however, there is a significant difference in their corresponding energy levels, which indicates that they are two distinct states. By comparing Figs. 6(a) and 7(a), it can be seen that as the value of $D$ increases, the “backbending” effect in $\beta(E)$ intensifies. As a result, the positions of the third-order independent transition and the phase transition point to become increasingly close to each other. This phenomenon is also observed in the Potts model, suggesting that it is characteristic of first-order phase transitions. The “backbending” phenomenon is a characteristic signal of phase coexistence in first-order phase transitions. When the primary phase transition of the system is of the first order, the maximum error between the positions of the third-order independent transition obtained by both methods is 0.002, considered negligible.
@@
In Fig. 9, we present the temperature-dependent variations of the average perimeter and its first derivative across different parameters. Our primary focus is to analyze the behavior of the first derivative. In the figure, red vertical lines mark the positions of the critical points, while purple lines indicate the locations of the identified third-order dependent transitions. Corresponding temperature values are annotated next to these lines in matching colors. When compared with the results obtained from microcanonical inflection point analysis, we find that the discrepancies between the two methods are approximately 0.03. Additionally, in Fig. 9(g), we observe a notable distinction, where a weak signal of a third-order dependent transition is detected and marked by a purple vertical line. Through careful analysis of this signal, we observed that the local minimum shifts upward as the system size increases. Consequently, we infer that, in the limit of infinite system size, this signal vanishes, aligning with the results from our microcanonical analysis. In Fig. 9(h), for $D=1.97$, where the system’s primary phase transition is of the first order, it is observed that this signal completely disappears. This phenomenon further implies that third-order dependent transition signals may not exist when the system’s primary phase transition is of the first order.
Figure 10: (Distribution maps of isolated spins and the largest cluster). This figure illustrates the spatial distribution of isolated spins and the largest cluster at different temperatures for various $D$ values, derived from simulations on a $200\times 200$ system. The largest cluster is shown in gray, while all other clusters are displayed uniformly in white. Isolated spins are represented by dots, with blue, green, and red dots corresponding to the $-1$, 00, and $+1$ spin states, respectively.
We thoroughly analyzed the distribution of isolated spins, the largest cluster, and the overall cluster structure to elucidate the characteristics of the third-order transition. Figures 10(a)-(d) depict the distribution of isolated spins and the largest cluster when the primary phase transition of the system is of the second order ($D=1.0$). In these figures, the largest cluster is represented in gray, while all other clusters are consistently shown in white. Isolated spins are shown as dots, with blue, green, and red dots representing isolated spins in the $-1$, 00, and $+1$ spin states, respectively. The figures reveal that the number of isolated spins peaks near the third-order independent transition point. Beyond this point, the largest cluster becomes increasingly fragmented by isolated spins, which subsequently coalesce into smaller clusters, resulting in a decline in their number. Figures 10(e)-(h) depict the distribution of isolated spins and the largest cluster when the primary phase transition of the system is of the first order ($D=1.97$). We observe that isolated spins in the 00 state are dominant before the third-order independent transition, while isolated spins in the $-1$ and $+1$ states emerge as predominant after the transition. These figures offer ample evidence supporting our explanation of the abrupt increase in the number of isolated spins as observed in Figures 8(c)-(d).
Figure 11: (Cluster distribution map). This figure illustrates the spatial distribution of clusters at different temperatures for various $D$ values, based on simulations of a $200\times 200$ system. The largest cluster is shown in gray, while clusters in the $-1$, 00, and $+1$ spin states are depicted in blue, green, and red, respectively.
Figure 11 illustrates the cluster distribution at various $D$ values and temperatures $T$. In the figure, the largest cluster is shown in gray, while clusters in the $-1$, 00, and $+1$ states are depicted in blue, green, and red, respectively. Figures 11(a)-(d) display the cluster distribution when the primary phase transition of the system is of the second order ($D=1.0$). As the temperature increases, the largest cluster gradually breaks down into smaller clusters that are still larger than typical small clusters, before further fragmenting into even smaller clusters. This process progresses continuously, manifesting macroscopically as a continuous phase transition, accompanied by the emergence of a third-order dependent transition. Figures 11(e)-(f) illustrate the cluster distribution when the primary phase transition of the system is of the first order ($D=1.97$). The largest cluster also gradually “dissolves”; however, within the temperature range of our simulation, it persists throughout the system. However, near the phase transition point, as the temperature $T$ increases, large clusters undergo a sudden transition from the -1 state to the 0 state. This phenomenon is clearly analysed in Figures 11(g) and (h) and has been validated through Metropolis sampling, which confirms this behavior. In fact, it represents a phase transition from an ordered ferromagnetic state to an ordered paramagnetic state. Analysis of cluster changes during the cooling process reveals that smaller clusters gradually grow but fail to form significantly larger clusters. We hypothesize that during cooling, smaller clusters continuously flip and may transition into states identical to the largest cluster, ultimately merging with it. This dynamic process inhibits the formation of larger cluster blocks, resulting in discontinuous changes in the order parameter. Therefore, when the primary phase transition of the system is of the first order, no third-order dependent transition signals are detected. Macroscopically, this phenomenon is marked by the discontinuity of the first-order phase transition. From these results, we find that the third-order transition is a phenomenon that emerges as the phase transition process.
Figure 12: (Cluster distribution before and after the fourth-order independent transition). This figure illustrates the spatial distribution of clusters at different temperatures for various $D$ values, based on simulations of a $200\times 200$ system. The largest cluster in the 0 state is shown in gray, while clusters in the $-1$, 00, and $+1$ spin states are depicted in blue, green, and red, respectively.
As previously noted, when the primary phase transition of the system is of the first order, it shifts from an ordered ferromagnetic state to an ordered paramagnetic transitions. We hypothesize that, with a further increase in temperature, a transition from the ordered paramagnetic state to the disordered paramagnetic state may occur. To investigate this phenomenon, we performed a more detailed analysis of the high-temperature region using the microcanonical inflection point analysis method. A fourth-order independent transition was observed at $D=1.965$, $1.969$, $1.97$, and $2.0$. In the inset of Figure 7 (d), the position of the fourth-order independent transition is indicated by a purple vertical line. Furthermore, the corresponding inset in Figure 7 (a) is provided to assist in extracting the relevant temperature information. This result was further validated by Metropolis sampling. In Figure 12, we present snapshots of the system before and after the fourth-order independent transition at $D=1.97$. We observed that the temperature at which the ordered paramagnetic phase transitions to the disordered paramagnetic phase is very close to the temperature of the fourth-order independent transition. This indicates that the fourth-order independent transition acts as a marker for the onset of complete disorder in the high-temperature region and can function as a warning indicator in high-temperature systems. The data were organized and presented in the form of a phase diagram in Figure 13. In the diagram, OFP denotes the ordered ferromagnetic phase, DPP denotes the disordered paramagnetic phase, and OPP denotes the ordered paramagnetic phase. Information regarding the phase transition points is obtained through microcanonical inflection point analysis. Near $D=1.9$, data for three specific points $D=1.965$, $1.969$, and $1.97$ detail the phase transition points. Second-order phase transitions are indicated by solid lines, first-order transitions by dashed lines, and the transition from the ordered paramagnetic phase to the disordered paramagnetic phase at high temperatures—when the primary phase transition is of the first order—is marked by a dotted line. When the primary phase transition of the system is of the first order, the system undergoes a transition from the ordered paramagnetic phase to the disordered paramagnetic phase through a fourth-order independent transition.
Figure 13: (Blume-Capel model phase map). This diagram expands on the previous phase diagram of the Blume-Capel model. OFP denotes the ordered ferromagnetic phase, DPP denotes the disordered paramagnetic phase, and OPP denotes the ordered paramagnetic phase. Second-order phase transitions are indicated by solid lines, first-order transitions by dashed lines, and the boundary for the transition from the ordered paramagnetic phase to the disordered paramagnetic phase when $D$ exceeds 1.965 is marked by a dotted line.
@@
Using two different methods, we identified signals of both third-order independent and dependent transitions within the second-order phase transition regime, while only third-order independent transition signals were detected in the first-order phase transition regime. Through microcanonical inflection point analysis and Metropolis sampling, we determined that when $D$ exceeds 1.965, the system transitions from the ordered ferromagnetic state to the ordered paramagnetic state, and eventually to the disordered paramagnetic state. The transition from the ordered paramagnetic state to the disordered paramagnetic state corresponds to fourth-order independent transition. We extend the phase diagram of the Blume-Capel model.
@@ 4 Summary
In this paper, we employ the Wang-Landau (WL) sampling method to obtain the density of states (DOS) for the Blume-Capel model, while using the Metropolis sampling method to simulate the spin configurations and extract information on isolated spins and the average cluster perimeter. We performed both canonical and microcanonical ensemble analyses on the data extracted from these two methods, successfully determining the positions of the critical and first-order phase transition points. The results from both sampling methods showed only minor differences, providing strong evidence for the validity and accuracy of our findings.
Importantly, our study revealed two intriguing phenomena. One of these is that through microcanonical inflection point analysis, we observed that both third-order dependent and independent transitions coexist at the critical phase transition, while only third-order independent transitions were observed in the first-order phase transition. This observation was further verified through geometric analysis of the behavior of isolated spins and average cluster perimeters. Similar conclusions have been drawn in the Ising model [18] and the Potts model [19]. Through the analysis of isolated spin distributions, largest cluster distributions, and the evolution of clusters, we found that third-order transitions may represent a phenomenon that accompanies the occurrence of phase transitions. Building on these findings, we further hypothesize that third-order dependent transitions vanish in the presence of strong first-order phase transitions. Another finding is that in the high-temperature region of the first-order phase transition in the Blume-Capel model, we used microcanonical inflection point analysis to identify a fourth-order independent phase transition. Our Metropolis sampling analysis of the largest cluster of zero spins suggests that the temperature of the fourth-order independent transition aligns with the transition temperature from the ordered paramagnetic phase to the disordered paramagnetic phase. This discovery further enriches the existing phase diagram of the Blume-Capel model. These two discoveries offer significant theoretical support for our research on critical early-warning systems and provide valuable guidance for future studies on phase transitions.
This method has provided us with more comprehensive information about the system, while the identified third-order transitions form a crucial basis for our early warning efforts. We intend to utilize this method for further exploration.
@@ 5 Acknowledgments
We would to thank Professor Youjin Deng for his valuable discussions. This work is supported by the National Natural Science Foundation of China No.12304257.